# The number of cultural traits, evolving genealogies, and the descendant process

### v1

@@ 3 Results
One of our major aims is to study the dynamics and the stationary distribution of $(C_{n}(t))_{t\geq 0}$. The typed ALG is the appropriate genealogical structure for this. Since it results from the untyped one by superposition of a Poisson point process that lays down innovations at rate $\mu$ on every line, and every innovation on the tree is passed on to the sample by construction, the crucial quantity is the tree length $L_{n}$ of (10). (Indeed, since we consider $t\to\infty$, the relevant tree length is $\lim_{\tau\to\infty}L_{n}(\tau)$.) By the genealogical picture,
@@
that is, $C_{n}$ follows the Poisson distribution with random parameter $\mu L_{n}$. Here and in what follows (and in line with the standard monographs (Durrett, 2008; Wakeley, 2009) in mathematical population genetics and coalescent theory), we let the genealogical picture speak for itself. Our first goal now is to characterise the distribution of $C_{n}$ by calculating moments of $C_{n}$ of any order.
3.1 Calculation of moments
3.1.1 Number of cultural traits
The $m$-th factorial moment of a random variable $W\sim\mathrm{Poi}(\nu)$ is $\mathbb{E}\left[W^{\underline{m}}\right]=\nu^{m}$; this is due to the simple fact that the probability-generating function of $W$ is $g(z)=e^{\nu(z-1)}$, and $\mathbb{E}\left[W^{\underline{m}}\right]=g^{(m)}(1)$. By (15), the $m$-th factorial moment of $C_{n}$ thus becomes
Together with the identity
[…]
where {${}{m}{i}$ is a Stirling number of the second kind, this gives the moments of $C_{n}$ in terms of the moments of $L_{n}$ as
[…]
For example, the first four moments read
In particular, the variance is
The latter can be seen as an instance of the standard decomposition of the variance: $\mathbb{V}\left[C_{n}\right]=\mathbb{E}\left[\mathbb{V}\left[C_{n}\mid L_{n}\right]\right]+\mathbb{V}\left[\mathbb{E}\left[C_{n}\mid L_{n}\right]\right]$. So the first term, $\mathbb{E}\left[\mathbb{V}\left[C_{n}\mid L_{n}\right]\right]=\mu\mathbb{E}\left[L_{n}\right]=\mathbb{E}\left[C_{n}\right]$, contains the variability due to the innovation process, whereas the second term, $\mathbb{V}\left[\mathbb{E}\left[C_{n}\mid L_{n}\right]\right]=\mu^{2}\mathbb{V}\left[L_{n}\right]$, comes from the variance of the length of the genealogy.
@@ 3.1.2 Path integrals of birth-death processes
Now that we know the moments of $C_{n}$ in terms of the moments of $L_{n}$, we set out to determine the latter. Note that $L_{n}$ in (10) is a special case of the path integral of a function $f:\mathbb{N}_{0}\to\mathbb{R}_{\geq 0}$ over a general continuous-time birth-death process $(Z(t))_{t\geqslant 0}$ on $\mathbb{N}_{0}$ with unique absorbing state 0, namely
@@
We assume throughout [^3] Note that, on this finite state space, we need not assume that $f$ is nondecreasing for large arguments, as required for $(Z(t))_{t\geqslant 0}$ on $\mathbb{N}_{\geq 0}$ in Stefanov and Wang (2000). that $f(0)=0$ and $f(i)>0$ for $i\in[N]$; in particular, this implies that the integral has no contribution from beyond the time of absorption in 0. The mathematical features of $\mathcal{I}_{i}(f)$, especially its expectation and its higher-order moments, have been well studied (Puri, 1966, 1968; Jerwood, 1970; McNeil, 1970; Goel and Richter-Dyn, 1974; Norden, 1982; Hernandez-Suarez and Castillo-Chavez, 1999; Stefanov and Wang, 2000; Pollett and Stefanov, 2002; Pollett, 2003; Crawford et al., 2018; Hobolth et al., 2019). Here we restrict $(Z(t))_{t\geqslant 0}$ to $[N]_{0}$. In Appendix B, we outline how to derive the moments of the path integral via first-step analysis; below we summarise the results.
@@
where the empty product is 1 (p. 162/163 in Stefanov and Wang (2000); note that, for $\ell\leq j$, $\eta_{\ell j}$ equals their $H_{\ell}(j)/\mu_{\ell}$; see also Hernandez-Suarez and Castillo-Chavez (1999) for the special case of $i=1$, and see eq. (1) of Stefanov and Wang (2000) for the limit $N\to\infty$). Here, $\eta_{\ell j}$ represents the expected total sojourn time in state $j$ measured from the moment where the process reaches state $\ell$ for the first time until the moment where it reaches state $\ell-1$ for the first time; for lack of reference and convenience of the reader, we prove this fact in Appendix A. Likewise, $\zeta_{ij}$ is the expected total sojourn time of $(Z(t))_{t\geqslant 0}$ in state $j$ before the process is absorbed in 0, given it started in $i$ (Stefanov and Wang, 2000, Eq. (4)). Generalising (21), the $m$-th moment of the path integral is given by
@@
3.1.3 Tree length
Let us now apply the results above to our cultural evolution model. With the choice $(Z(t))_{t\geqslant 0}=(Y(\tau))_{\tau\geqslant 0}$, $f=\mathrm{id}$, and $i=n$, we have $L_{n}=\mathcal{I}_{n}(f)$. With the transition rates of $(Y(\tau))_{\tau\geqslant 0}$ in (9), (23) evaluates to
@@ 3.1.3 Tree length
Inserting these results into eqs.(18) and (19), we obtain the moments of $C_{n}$ in a closed form, see Figure 4 for expectation and variance. This also illustrates the (smooth) transition from the subcritical phase with a small number of traits to the supercritical phase, where the mean number of traits increases steeply with $s$, as previously observed in similar models (Nakamura et al., 2020; Kobayashi et al., 2018; Kobayashi et al., 2021). Comparing with the variance decomposition below (19), the figure also shows that, for $s>1$, the variability due to innovations, which equals $\mathbb{E}\left[C_{N}\right]$, is small relative to $\mathbb{V}\left[C_{N}\right]$, so the major contribution to $\mathbb{V}\left[C_{N}\right]$ is $\mu^{2}\mathbb{V}\left[L_{N}\right]$, which stems from the variance of the tree length. We compare the result to simulations in the next section.
Figure 4: $\mathbb{E}\left[C_{N}\right]$ (solid) and $\mathbb{V}\left[C_{N}\right]$ (broken) as functions of $s$. $N=100,\mu=0.1,u=1$.
3.2 Forward simulations
3.2.1 Procedure
We carried out individual-based simulations of the forward process $(\Phi(t))_{t\in[0,t_{\rm{max}}]}$ described in Section 2.2. Recall that death, learning, and innovation events occur at total rates $Nu$, $Ns$, and $N\mu$, respectively. We work with $u=1$ throughout, so that a time interval of length 1 corresponds to the expected life time of an individual, that is, one generation. The waiting time and kind for each next event is determined by drawing exponential random numbers with parameter $N(u+s+\mu)$ and then letting a death, a learning, and an innovation event happen with probability $Nu/(N(u+s+\mu))$, $Ns/(N(u+s+\mu))$, and $(N\mu/N(u+s+\mu))$, respectively. When a death event occurs, a uniformly-chosen individual loses all traits. When learning occurs, we choose two individuals uniformly and with replacement, namely a teacher (that is, a cultural parent) and a learner, and all traits of the teacher are added to the learner’s repertoire. When an innovation occurs, a uniformly-chosen individual acquires a new trait that has never existed before. Within the limited memory of the computer, this is feasible by "forgetting" the extinct traits and reusing the corresponding storage.
We start with $k_{i}(0)=\varnothing$ for every $i\in[N]$ and compute realisations $c_{n}(t)$ of $C_{n}(t)$ at $t=0,1,...,t_{\rm{max}}$. We take $t_{\rm{max}}$ large enough so that the effect of the initial state can be neglected and denote by $[t_{\rm{max}}]_{0}$ the set $\{0,1,\ldots,t_{\rm{max}}\}$. Note that, by definition (see (5)), $C_{n+1}(t)\geq C_{n}(t)$ for all $t$.
3.2.2 Time series
A cutout of the time series $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ of a subcritical case ($s<u$) and a supercritical case ($s>u$) are shown in Figure 5. More detail is to be seen in Figure 6, which shows the time series $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ and $(c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$ for the last 400 generations in both cases. In the subcritical case, both $(c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$ and $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ fluctuate moderately around some constant value in a more or less symmetric manner. In contrast, we find a characteristic sawtooth behaviour in the supercritical case: $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ tends to increase linearly (at roughly constant slope in every sawtooth) for some period, followed by a rapid collapse to a small value at a random time. We will discuss the sawtooth behaviour later. For now, note that $t_{\rm{max}}$ is large enough for the mass extinction of traits to occur sufficiently often during the simulated period. Despite the dynamics, we are thus in the stationary situation. Let us therefore first consider the moments of (the stationary) $C_{n}$ and then turn to the dynamics.
Figure 5: Time series $\big{(}c_{N}(t)\big{)}_{t\in[t_{\rm{max}}]_{0}}$ in the subcritical case (upper panel, $s=0.9$) and the supercritical case (lower panel, $s=1.3$). The last $10^{4}$ generations of a long run with $t_{\rm{max}}=10^{5}$ are shown. $N=100,\mu=0.1,u=1.$
Figure 6: Time series $\big{(}c_{1}(t)\big{)}_{t\in[t_{\rm{max}}]_{0}}$ (green) and $\big{(}c_{N}(t)\big{)}_{t\in[t_{\rm{max}}]_{0}}$ (blue) for the last 400 generations in a long run with $t_{\rm{max}}=10^{5}$ in the subcritical case (upper panel, $s=0.9$) and the supercritical case (lower panel, $s=1.3$). $N=100,\mu=0.1,u=1$.
@@ 3.2.3 Moments of $C_{n}$
As a consistency check, we calculate the empirical time average and the corresponding variance, that is,
$\bar{c}_{n}\mathrel{\mathop{:}}=\frac{1}{t_{\rm{max}}+1}\sum_{i=0}^{t_{\rm{max}}}c_{n}(t)$ (29)
and
$s_{c}^{2}\mathrel{\mathop{:}}=\frac{1}{t_{\rm{max}}}\sum_{i=0}^{t_{\rm{max}}}\big{(}c_{n}(t)-\bar{c}_{n}\big{)}^{2}$ (30)
@@
for various values of $n$ and compare $\bar{c}_{n}$ and $\sqrt{s_{c}^{2}}$ with $\mathbb{E}\left[C_{n}\right]$ and $\sqrt{\mathbb{V}\left[C_{n}\right]}$ according to eqs.(18) and (19). Figure 7 shows that they are in good agreement in both the sub- and supercritical case. The most important difference is that, in the supercritical case, $\mathbb{E}\left[C_{n}\right]$ saturates for relatively small $n$; that is, almost all traits of the population are already contained in a small sample, as one may already suspect from Figure 6, and as will be analysed in what follows. Of course, in addition, $\mathbb{E}\left[C_{n}\right]$ and $\mathbb{V}\left[C_{n}\right]$ are altogether substantially higher in the supercritical case; also compare Figure 4, which shows the dependence on $s$. Note also that, in the supercritical case, $\mathbb{E}\left[C_{n}\right]\ll\mathbb{V}\left[C_{n}\right]$ for all $n$, so the variance of the tree length dominates the variability due to innovations for all $n\in[N]$, as already discussed in the context of Figure 4 for $n=N$.
Figure 7: Analytic and simulation results of mean and standard deviation of $C_{n}$ as a function of $n$ in the subcritical case (upper panel, $s=0.9$) and the supercritical case (lower panel, $s=1.3$). Solid and broken lines show $\mathbb{E}\left[C_{n}\right]$ and $\sqrt{\mathbb{V}\left[C_{n}\right]}$ of (18) and (19), respectively. Large and small discs represent $\bar{c}_{n}$ and $\sqrt{s_{c}^{2}}$ of (29) and (30). $N=100,\mu=0.1,u=1,t_{\rm{max}}=10^{5}$.
3.2.4 Marginal and joint distributions of $C_{n}$
Figure 8 shows the histograms of $(\log_{10}c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$ and $(\log_{10}c_{N})_{t\in[t_{\rm{max}}]_{0}}$ in the sub- and supercritical cases. In contrast to the subcritical case, the histogram of $(\log_{10}c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$ in the supercritical case is bimodal. It peaks sharply at 0, where most of the mass is located, decays quickly from there, and has a second peak around $c_{1}=10^{2.3}$. So, an individual is ignorant most of the time, but occasionally becomes very knowledgeable. The histogram of $(\log_{10}c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ has a single peak near the second peak of the histogram of $(\log_{10}c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$. What this means becomes clear in the heat maps (Figure 9), which show how the $c_{n}$ covary. In the subcritical case, the correlation between $c_{1}$ and $c_{N}$ is altogether weak; $c_{1}$ is bimodal in the rare case that $c_{N}$ is large. A similar tendency exists between $c_{2}$ and $c_{N}$. In contrast, the correlation is very strong in the supercritical case. When $c_{N}$ is not too small, the distribution of $c_{1}$ is bimodal, meaning that there are two extreme types of individuals: those who know nearly nothing and those who have almost all traits currently present in the population, with very few intermediate cases. An analogous observation applies to $c_{2}$.
Figure 8: Histograms of $\big{(}\log_{10}c_{1}(t)\big{)}_{t\in[t_{\rm{max}}]_{0}}$ (green) and $\big{(}\log_{10}c_{N}(t)\big{)}_{t\in[t_{\rm{max}}]_{0}}$ (violet) obtained by simulation runs of the subcritical case (upper panel, $s=0.9$) and the supercritical case (lower panel, $s=1.3$); the overlap area of the two histograms appears blue-green. $c_{1}=0$ was observed 76376 times (upper panel) and 64394 times (lower panel), which are not shown. $N=100,\mu=0.1,u=1,t_{\rm{max}}=10^{5}$.
Figure 9: Heat maps for $(c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$ and $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ and for $(c_{2}(t))_{t\in[t_{\rm{max}}]_{0}}$ and $(c_{N}(t))_{t\in[t_{\rm{max}}]_{0}}$ in the subcritical case (upper panel, $s=0.9$) and the supercritical case (lower panel, $s=1.3$) in a long run with $t_{\rm{max}}=10^{5}$. The angle bisector represents $c_{1}(t)=c_{N}(t)$ for all $t$, that is, individual 1 has all traits present in the population at time $t$. $N=100,\mu=0.1,u=1$.
3.3 Genealogies and backward simulations
3.3.1 Distributions of tree length, $L_{n}$
To understand the behaviour observed in the forward simulations, we now recall that $C_{n}\sim\text{Poi}(\mu L_{n})$ together with the observation that, in the supercritical case, the variance of $C_{n}$ mainly stems from the variance of $L_{n}$. We therefore expect that the underlying genealogies play the decisive role and so turn to them now. Figure 10 shows the histograms of $\log_{10}\ell_{1}$ and $\log_{10}\ell_{N}$ (where $\ell_{n}$ is the realisation of the (stationary) $L_{n}$) obtained via (10) by simulating the untyped ALG $(\Lambda(\tau))_{\tau\geq 0}$ according to (6b) until extinction, a large number of times. (Since we need the full untyped ALG in what follows, we refrained from simulating the more efficient line-counting process $(Y_{n}(\tau))_{\tau\geq 0}$ instead.) Note that the left peaks of the histograms of $\log_{10}\ell_{1}$ translate into the left peaks of the histograms of $(\log_{10}c_{1}(t))_{t\in[t_{\rm{max}}]_{0}}$ in that short tree lengths mean high chances for no or only a small number of innovation events, resulting in (almost) ignorant individuals; and vice versa for the other peaks.

### v2

@@ 3 Results
The typed ALG is the appropriate genealogical structure to study the dynamics and the stationary distribution of $(C_{n}(t))_{t\geq 0}$. Since the typed ALG results from the untyped one by superposition of a Poisson point process that lays down innovations at rate $\mu$ on every line, and every innovation on the genealogy is passed on to the sample, the crucial quantity is the length of the genealogy, $L_{n}$ of (10). By the genealogical picture,
@@
that is, $C_{n}$ is a mixed Poisson random variable: a Poisson random variable with a random parameter $\mu L_{n}$. So the genealogies, and $L_{n}$ in particular, are the fundamental quantities, which are also of independent interest, and $C_{n}$ contains additional randomness due to the Poissonisation.
We will, in what follows, consider the moments and the distributions of $L_{n}$ and $C_{n}$ and then turn to the dynamics of $(L^{t}_{n})_{t\geq 0}$ and $(C_{n}(t))_{t\geq 0}$, throughout with an eye on the additional fluctuations induced by the innovation process.
3.1 The moments
@@ 3.1.1 Path integrals of birth-death processes
We start by observing that $L_{n}$ in (10) is a special case of the path integral of a function $f:\mathbb{N}_{0}\to\mathbb{R}_{\geq 0}$ over a general continuous-time birth-death process $(Z(t))_{t\geqslant 0}$ on $\mathbb{N}_{0}$ with unique absorbing state 0, namely
@@
We assume throughout [^3] Note that, on this finite state space, we need not assume that $f$ is nondecreasing for large arguments, as required for $(Z(t))_{t\geqslant 0}$ on $\mathbb{N}_{\geq 0}$ in Stefanov and Wang (2000). that $f(0)=0$;
in particular, this implies that the integral has no contribution from beyond the time of absorption in 0. The mathematical features of $\mathcal{I}_{i}(f)$, especially its expectation and higher-order moments, have been well studied (Puri, 1966, 1968; Jerwood, 1970; McNeil, 1970; Goel and Richter-Dyn, 1974; Norden, 1982; Hernandez-Suarez and Castillo-Chavez, 1999; Stefanov and Wang, 2000; Pollett and Stefanov, 2002; Pollett, 2003; Crawford et al., 2018; Hobolth et al., 2019). Here we restrict $(Z(t))_{t\geqslant 0}$ to $N]_{0}$. In Appendix B, we outline how to derive the moments of the path integral via first-step analysis; below we summarise the results.
@@
where the empty product is 1 (p. 162/163 in Stefanov and Wang (2000); note that, for $\ell\leq j$, $\eta_{\ell j}$ equals their $H_{\ell}(j)/\mu_{\ell}$; see also Hernandez-Suarez and Castillo-Chavez (1999) for the special case of $i=1$, and see eq. (1) of Stefanov and Wang (2000) for the special case $N=\infty$). Here, $\eta_{\ell j}$ represents the expected total sojourn time in $j$ measured from the moment where the process reaches state $\ell$ for the first time until the moment where it reaches state $\ell-1$ for the first time; for lack of reference and convenience of the reader, we prove this fact in Appendix A. Consequently, $\zeta_{ij}$ is the expected total sojourn time of $(Z(t))_{t\geqslant 0}$ in $j$ before the process is absorbed in 0, given it started in $i$ (Stefanov and Wang, 2000, Eq. (4)). Generalising (17), the $m$-th moment of the path integral is given by
@@
3.1.2 Moments of $L_{n}$
Let us now apply the results above to our genealogies. With the choice $(Z(t))_{t\geqslant 0}=(Y(\tau))_{\tau\geqslant 0}$, $f=\mathrm{id}$, and $i=n$, we have $L_{n}=\mathcal{I}_{n}(f)$. With the transition rates of $(Y(\tau))_{\tau\geqslant 0}$ in (9), (19) evaluates to
@@ 3.1.3 Moments of $C_{n}$
It is now an easy exercise to translate the moments of $L_{n}$ into the moments of $C_{n}$. The $m$th factorial moment of a random variable $W\sim\mathrm{Poi}(\nu)$ is $\mathbb{E}\left[W^{\underline{m}}\right]=\nu^{m}$; this is due to the simple fact that the probability generating function of $W$ is $g(z)=e^{\nu(z-1)}$, and $\mathbb{E}\left[W^{\underline{m}}\right]=g^{(m)}(1)$. By (15), the $m$th factorial moment of $C_{n}$ thus becomes
Together with the identity
[…]
For example, the first four moments read
In particular, the variance is
The latter can be seen as an instance of the standard decomposition of the variance: $\mathbb{V}\left[C_{n}\right]=\mathbb{E}\left[\mathbb{V}\left[C_{n}\mid L_{n}\right]\right]+\mathbb{V}\left[\mathbb{E}\left[C_{n}\mid L_{n}\right]\right]$. So the first term, $\mathbb{E}\left[\mathbb{V}\left[C_{n}\mid L_{n}\right]\right]=\mu\mathbb{E}\left[L_{n}\right]=\mathbb{E}\left[C_{n}\right]$, contains the variability due to the innovation process, whereas the second term, $\mathbb{V}\left[\mathbb{E}\left[C_{n}\mid L_{n}\right]\right]=\mu^{2}\mathbb{V}\left[L_{n}\right]$, comes from the variance of the length of the genealogy. Inserting (23) and (24) into (27) and (28), we obtain the first two moments of $C_{n}$ in closed form.
For numerical and simulation results, we work with $u=1$ throughout, so one time unit corresponds to the expected life time of an individual, that is, one generation. Figure 4 shows the two components of $\mathbb{V}\left[C_{N}\right]$ and their ratio as functions of $s$. First, the solid line $\mu\mathbb{E}\left[L_{N}\right]=\mathbb{E}[C_{N}]$ illustrates the (smooth) transition from the subcritical phase with a small number of traits to the supercritical phase, where the mean number of traits increases steeply with $s$, as previously observed in similar models (Nakamura et al., 2020; Kobayashi et al., 2018; Kobayashi et al., 2021). According to the variance decomposition above, $\mu\mathbb{E}\left[L_{N}\right]$ is, at the same time, the variability of $C_{N}$ due to innovations. The figure shows that, for small values of $s$, it is the (by far) dominating contribution to $\mathbb{V}[C_{N}]$; but the proportion decreases with $s$ and, for $s>1$, $\mu^{2}\mathbb{V}[L_{N}]$, the variability of $C_{N}$ due to the length of the genealogy, is the major contribution to $\mathbb{V}[C_{N}]$. The picture is very similar for the variance components of $C_{1}$ (not shown).
Figure 4: The variance components $\mu\mathbb{E}\left[L_{N}\right]$ (solid, left scale) and $\mu^{2}\mathbb{V}\left[L_{N}\right]$ (dashed, left scale), along with the proportion of the former, $\mu\mathbb{E}\left[L_{N}\right]/\mathbb{V}\left[C_{N}\right]$ (dot-dashed, right scale) as functions of $s$. $N=100,\mu=0.1,u=1$.
@@
Figure 5 shows $\mathbb{E}\left[C_{n}\right]$ and $\sqrt{\mathbb{V}\left[C_{n}\right]}$ for fixed $s$ as functions of $n$ in both the sub- and the supercritical case. The most important difference is that, in the supercritical case, $\mathbb{E}\left[C_{n}\right]$ saturates for relatively small $n$; that is, almost all traits of the population are already contained in a small sample, a fact that will turn out as crucial in the next section. Of course, in addition, $\mathbb{E}\left[C_{n}\right]$ and $\sqrt{\mathbb{V}\left[C_{n}\right]}$ are altogether substantially higher in the supercritical case, as we have already seen for $n=N$ in Figure 4.
Figure 5: $\mathbb{E}\left[C_{n}\right]$ (solid) and $\sqrt{\mathbb{V}\left[C_{n}\right]}$ (broken) as functions of $n$ in the subcritical case ($s=0.9$, upper panel) and the supercritical case ($s=1.3$, lower panel). The graphs of $\mathbb{E}\left[L_{n}\right]$ and $\sqrt{\mathbb{V}\left[L_{n}\right]}$ are not shown since they are almost indistinguishable from those of $\mathbb{E}\left[C_{n}\right]$ and $\sqrt{\mathbb{V}\left[C_{n}\right]}$ except for the scaling factor of $\mu$. $N=100,u=1,\mu=0.1$.
3.2 Marginal and joint distributions of $L_{1}$ and $L_{N}$, and of $C_{1}$ and $C_{N}$
Figure 6 shows the histograms of $\log_{10}\ell_{1}$ and $\log_{10}\ell_{N}$ (where $\ell_{n}$ is the realisation of $L_{n}$) obtained via (10) by simulating the untyped ALG $(\Lambda(\tau))_{\tau\geq 0}$ following (6b) until extinction, a large number of times.